# Distilling Textual Priors from LLM to Efficient Image Fusion

### v2

@@ 4.2 Comparison on IVF Tasks
Table 2: Quantitative comparison on Harvard Medical Dataset. The number in () is non-trainable parameters but used in inference period. Bold and underlined values indicate the best and second-best results respectively.
Method Params(M) PET-MRI CT-MRI SPECT-MRI
[…]
PSLPT 3.06 0.815 0.548 0.373 0.810 0.502 0.432 0.933 0.359 0.325
EMFusion 0.78 1.221 0.685 0.783 1.266 0.552 0.475 1.212 0.665 0.692
MSRPAN 0.39 1.182 0.581 0.799 1.261 0.436 0.455 1.153 0.525 0.560
SwinFusion 13.04 0.725 0.703 0.683 0.579 0.522 0.545 0.684 0.744 0.720
Zero 19.1 1.162 0.635 0.774 1.199 0.320 0.582 1.180 0.582 0.681
U2Fusion 0.63 0.494 0.460 0.292 0.042 0.074 0.489 0.479 0.419 0.696
CDDFuse 2.4 1.227 0.650 0.765 1.224 0.526 0.530 1.169 0.786 0.719
Text-IF 63.8(+151) 1.232 0.640 0.690 1.313 0.542 0.561 1.200 0.747 0.715
Ours-teacher 40.3(+151) 1.223 0.909 0.782 1.313 0.641 0.657 1.210 0.888 0.746
Ours-distilled 4.1 1.243 0.929 0.784 1.312 0.653 0.657 1.202 0.887 0.747
4.3 Comparison on Medical Datasets
To demonstrate the generalization capability of our method, we evaluate its performance on medical image fusion tasks. As shown in Table 2, both our teacher and student networks achieve superior performance across different modality pairs. The improvement is particularly significant in structural preservation metrics (SSIM and $Q^{AB/F}$), which is crucial for medical applications.
Figure 4: Compare with previous SOTA text-guided image fusion method Text-IF on Havard Medical Image Fusion Datasets. More can be found in appendix.
For SPECT-MRI fusion, our method shows notable advantages in preserving functional information while maintaining anatomical details. In CT-MRI fusion, where structural alignment is critical, our approach achieves the highest SSIM scores, indicating better structural preservation. The PET-MRI results further confirm our method’s effectiveness in handling multi-modal medical images with different characteristics. Figure4 presents visual examples from different medical modalities. Our fusion results exhibit better detail preservation and contrast enhancement, which is essential for clinical applications. The distilled student network maintains these advantages while significantly reducing computational requirements, making it more practical for clinical deployment.
4.4 Ablation Study
Figure 5: Overview of our ablation study
4.4.1 Analysis of Teacher Network
Effect of Text Guidance.
To validate the effectiveness of text guidance, we compare our full model with its variant without text information. As shown in Table 3, removing text guidance leads to a significant performance drop across all metrics (EN: -0.023, VIF: -0.083, $Q^{AB/F}$: -0.05). This demonstrates that text descriptions provide valuable semantic information that helps the model better understand and fuse image features. This also demonstrates the effectiveness of our model distillation. The text guidance acts as an advanced form of reference information, offering rich contextual information beyond pixel-level features.
Table 3: Ablation study on text guidance. Results on MSRS test set.
[…]
Non-text 6.726 0.977 0.682
With text 6.749 1.060 0.732
Distilled 6.763 1.075 0.734
Impact of Different Loss Components.
We evaluate the contribution of each loss component by systematically removing them from the full model. As shown in Table 4, all loss components positively impact the final performance, although their importance varies. The intensity loss ($\mathcal{L}_{int}$) and gradient loss ($\mathcal{L}_{grad}$) have the most significant influence, highlighting their essential role in preserving structural information. The SSIM loss ($\mathcal{L}_{ssim}$) aids in maintaining perceptual quality, while the color loss ($\mathcal{L}_{color}$) ensures the natural appearance of fused results.
Table 4: Ablation study on loss components for the teacher network. Results are reported on the MSRS test set.
[…]
$\mathcal{L}_{int}$ 6.695 1.027 0.697
$\mathcal{L}_{int}+\mathcal{L}_{color}$ 6.696 1.025 0.698
$\mathcal{L}_{int}+\mathcal{L}_{grad}+\mathcal{L}_{ssim}$ 6.731 1.058 0.730
$\mathcal{L}_{int}+\mathcal{L}_{color}+\mathcal{L}_{grad}$ 6.733 1.059 0.733
Feature Fusion Methods.
We investigate different fusion embedding strategies to combine features from infrared and visible images. The comparison includes concatenation, addition, and our proposed adaptive fusion method. Results demonstrate that our adaptive fusion approach achieves superior performance by dynamically adjusting the fusion weights based on spatial and channel features.
Table 5: Comparison of different fusion modules. Results on MSRS test set.
[…]
Cat & 1×1 conv 6.738 1.059 0.728
Unlearn weight atten 6.667 0.971 0.655
Dynamic weight conv 6.737 1.058 0.729
Multiscale conv 6.743 1.053 0.720
Ours 6.749 1.060 0.732
4.4.2 Analysis of Student Network
Effect of Model Size.
A notable finding in our distillation experiments is that reducing the model size by up to 90% (from 40.3M to 4.1M parameters if not calculate CLIP) does not significantly impact performance. As shown in Table 6, the compact student model maintains comparable or even slightly better results across different metrics. This suggests that the original model might be over-parameterized for the current fusion task, and the knowledge distillation process effectively transfers essential fusion capabilities to a much smaller architecture.
Table 6: Impact of model size on distilled network. Results on MSRS test set.
[…]
48dim 39.1 6.780 1.071 0.734
32dim 15.8 6.768 1.071 0.734
16dim 4.1 6.763 1.075 0.734
Impact of Distillation Losses.
We evaluate the contribution of different components in our distillation framework. The results in Table 7 show that result-level supervision (matching the teacher’s output) provides strong guidance, significantly improving the student model’s performance. The feature distillation loss, on the other hand, plays a key role in preserving feature-level consistency, ensuring that the student model captures important intermediate representations from the teacher. By introducing feature distillation, the model learns high-frequency details and intermediate features, enriching the information. This improves metrics like $EN$ and $Q^{AB/F}$, though it may slightly reduce the $VIF$ score due to imperfect alignment with perceptual fidelity.
Table 7: Ablation study on distillation losses (MSRS dataset).
[…]
$\mathcal{L}_{base}$ 6.728 0.989 0.681
$\mathcal{L}_{base}+\mathcal{L}_{res}$ 6.761 1.082 0.732
$\mathcal{L}_{base}+\mathcal{L}_{feat}+\mathcal{L}_{res}$ 6.763 1.075 0.734
Our ablation studies reveal several key insights:
5 Conclusion
We address the weak semantic understanding of traditional image fusion methods and the high cost of text-guided approaches by distilling textual priors and eliminating the need for text guidance during inference. Our approach, which leverages a teacher-student architecture and tailored prior distillation, significantly reduces the model size while retaining high performance. Extensive experiments and ablation studies validate the effectiveness of our method, highlighting its ability to achieve a strong trade-off between computational efficiency and fusion quality. We also analyze the contribution of individual components, offering valuable insights into their roles in the overall performance. Interestingly, we observe that the distilled model occasionally outperforms the teacher model, a phenomenon we discuss further in the appendix.
Appendix A Teacher Network Loss Terms
This section will introduce complete mathematical formulations for all loss terms which we used in the train process of teacher network.
$\mathcal{L}_{int}=\frac{1}{HW}\|I_{f}-\max(I_{vis}^{g},I_{ir}^{g})\|_{1}$ (17)
where $F_{CbCr}$ denotesthe transfer function of RGB to YCbCr.
These loss terms work together to ensure high-quality image fusion: $\mathcal{L}_{int}$ maintains the maximum intensity information from both source images, $\mathcal{L}_{SSIM}$ preserves structural details through the SSIM[26] metric, $\mathcal{L}_{grad}$ ensures edge consistency through gradient preservation, and $\mathcal{L}_{color}$ maintains natural color appearance. The text-dependent weights $\alpha(t)$ allow dynamic adjustment of each loss component’s contribution based on the specific fusion requirements described in the text prompt.
Appendix B Additional Quantitative Results
B.1 More Comparison on Infrared-Visible Tasks
Due to space limitations in the main paper, we present additional evaluation metrics here. Table 10 provides a comprehensive comparison across three datasets (MSRS, M3FD, and RoadScene) using five metrics: Entropy (EN[23]), Mutual Information (MI[42]), Spatial Frequency (SF[36]), Visual Information Fidelity (VIF[24]), and Quality of Gradient-based Fusion ($Q^{AB/F}$). Our method demonstrates consistent superiority across most metrics, particularly in structural preservation (VIF, $Q^{AB/F}$[25]) and information content (EN, MI).
The distilled student network achieves comparable or even better performance than the teacher network while using only 4.1M parameters, compared to the teacher’s 40.3M (+151M fixed parameters). This significant reduction in model size (approximately 90%) without performance degradation validates the effectiveness of our knowledge distillation approach. Notably, on the MSRS dataset, our distilled model improves upon the teacher network across all metrics, suggesting that the distillation process not only transfers knowledge but also helps regularize the model.
B.2 More Comparison on Medical Images
Table 11 extends our evaluation to medical image fusion tasks across three modality pairs (PET-MRI, CT-MRI, and SPECT-MRI) using additional metrics including SSIM, VIF, $Q^{AB/F}$, MI, and EN. The results demonstrate our method’s strong generalization capability across different medical imaging modalities.
For PET-MRI fusion, our distilled model achieves the highest scores in SSIM (1.243) and VIF (0.929), indicating superior structural preservation and perceptual quality. In CT-MRI fusion, both teacher and student networks maintain high performance, particularly in structural metrics (SSIM: 1.313) and edge preservation ($Q^{AB/F}$: 0.657). The SPECT-MRI results further validate our method’s effectiveness, with the distilled network showing exceptional performance in information preservation (MI: 4.371) while maintaining high structural fidelity.
Appendix C Extended qualitative Results
C.1 Visual Comparison on Infrared-Visible Images
Figure 8 presents additional visual comparisons across different challenging scenarios from the MSRS, M3FD, and RoadScene datasets. Our method demonstrates superior performance in: 1) Preserving thermal information from infrared images while maintaining natural visible details 2) Handling extreme lighting conditions and complex textures 3) Maintaining clear object boundaries and structural integrity 4) Achieving balanced fusion results without introducing artifacts
C.2 Visual Comparison on Medical Images
Figure 9 showcases detailed visual comparisons across different medical imaging modalities. The results highlight our method’s advantages in: 1) Preserving fine anatomical structures from both modalities 2) Maintaining clear tissue boundaries and contrast 3) Effectively combining functional information (PET/SPECT) with anatomical details (MRI) 4) Producing clinically meaningful fusion results with enhanced diagnostic value
Appendix D Extend Ablation Studies
D.1 Text Prior Analysis
To better understand how textual priors influence the fusion process, we conduct a detailed analysis of their effects at different network levels. As shown in Figure 6, we visualize the feature maps before and after applying text guidance, as well as the final fused results under different text descriptions.
The visualization reveals that text guidance significantly impacts the model’s attention to different image regions. For instance, when processing a low-contrast nighttime scene, the text prior helps the model better distinguish the vehicle’s contours from the background. This is evident in the feature maps, where the “After prompt guidance” visualization shows more defined object boundaries and clearer separation between foreground and background elements compared to the “Before prompt guidance” state.
Figure 6: Effect of text guidance on level 1 and fused image.
Interestingly, different text descriptions (e.g., “Low Contrast” vs. “Low Light”) lead to distinct feature emphasis patterns. Under the “Low Light” condition, the model pays more attention to illumination-related features, while the “Low Contrast” description prompts the model to focus more on edge and boundary information. This demonstrates the model’s ability to adapt its fusion strategy based on the provided textual context.
The “Distilled” results show that our student network successfully inherits these text-guided attention patterns without requiring explicit textual input during inference. This validates the effectiveness of our prior distillation approach in preserving the semantic understanding capabilities of the teacher network while significantly reducing computational overhead.
D.2 Knowledge Distillation Analysis
An intriguing observation from our experiments is that the distilled student network occasionally outperforms its teacher model. To investigate this phenomenon, we conducted a detailed analysis of the distillation process and its impact on model behavior.
Figure 7: When teacher perform extremely, the student learn more information.
Balance Between Structural and Color Information.
When training the teacher network with text guidance, we observed that increasing the weight of $\mathcal{L}_{ssim}$ leads to stronger edge and boundary preservation but may cause color degradation issues. Interestingly, during the distillation process, the student network appears to naturally balance these competing objectives. As shown in Figure 7, while the teacher model might exhibit color breakdown when heavily prioritizing structural information, the student network maintains edge clarity while simultaneously preserving natural color representation.
Impact of Teacher’s Behavior on Student Performance.
Our experiments reveal that the relationship between teacher and student performance is not strictly monotonic. We tested this by training teacher models with varying degrees of $\mathcal{L}_{ssim}$ emphasis. With an extreme weighting (10× the baseline), the teacher model becomes highly specialized in edge detection but sacrifices other aspects of image quality. Surprisingly, this extreme specialization enables the student to learn more comprehensive fusion strategies, potentially due to the clearer supervision signal for structural features.
However, such extreme settings are primarily valuable for experimental analysis rather than practical applications. In our final implementation, we adopt a moderately increased weight for $\mathcal{L}_{ssim}$ compared to our balanced version (where all loss components have equal weights). This moderate adjustment proves sufficient for the student network to achieve superior performance across multiple metrics while maintaining stable training dynamics.
This phenomenon might be attributed to several factors:
  • The distillation process acts as a form of regularization, preventing the student from inheriting potential overfitting behaviors from the teacher
  • The student network’s architecture may be better suited for balancing multiple objectives simultaneously
  • The knowledge distillation framework allows the student to learn from the teacher’s intermediate representations, potentially capturing complementary information that aids in overall fusion quality
D.3 Inference Time Analysis
To evaluate the computational efficiency, we analyze the inference times of different components across the previous state-of-the-art (SoTA) text-guided method Text-IF, our teacher network, and the distilled student network. For text generation, we use Qwen2-VL as the large language model (LLM), which operates on two NVIDIA RTX 4090 GPUs with the fast attention library[44]. Other timing experiments were conducted on a separate single NVIDIA RTX 4090 GPU, running with a batch size of 1 and an image resolution of $128\times 128$ pixels. Since these results may vary across different machines or even different states of the same machine, this comparison should be considered non-rigorous.
Table 8: Detailed inference time comparison (in milliseconds) for different components. LLM refers to the Qwen2-VL model used for text generation.
Method Data Load LLM CLIP Fusion Total
Text-IF 0.01 2261.2 115.7 152.1 2529.01
Teacher 0.01 2311.1 112.4 133.4 2556.91
Student 0.01 - - 46.1 46.11
As shown in Table 8, the LLM component dominates the computational cost, requiring over 2 seconds per image. CLIP encoding adds another 110-115ms overhead. While the teacher network’s fusion module takes 133.4ms, our distilled student network reduces this to just 46.1ms - a 65% reduction in fusion time. More importantly, by eliminating the need for LLM and CLIP during inference, our student network achieves a dramatic 98% reduction in total inference time (from 2556.91ms to 46.11ms).
D.4 Limitations
Table 9: Training time comparison for different epochs. All experiments were conducted on a single NVIDIA 4090 GPU.
Method 100 epochs 500 epochs
Text-IF 1.9h 11.1h
Teacher 2.0h 12.1h
Student (distill) 2.1h 13.5h
As shown in Table 9, our distillation approach requires slightly longer training time compared to both the baseline Text-IF and the teacher model. For a 100-epoch training cycle, the student network’s distillation process takes 2.1 hours, approximately 10.5% longer than Text-IF (1.9h) and 5% longer than the teacher model (2.0h). This pattern persists for longer training durations, with 500-epoch training requiring 13.5 hours for distillation compared to 11.1 hours for Text-IF and 12.1 hours for the teacher model.
Table 10: Quantitative comparison with SOTA methods on IVF tasks. Bold and underlined values indicate the best and second-best results respectively. Tra.P. = Trainable Parameters, Fix.P. = Fixed Parameters
@@ D.4 Limitations
Table 11: Quantitative comparison with SOTA methods on medical image fusion tasks. Bold and underlined values indicate the best and second-best results respectively. The additional parameters (in parentheses) are non-trainable but used during inference.
@@
Figure 8: More compare with different methods on different datasets.
Figure 9: More compare with different methods on different datasets.

### v3

@@ IV-B Comparison on IVF Tasks
TABLE I: Quantitative comparison with SOTA methods on IVF tasks across MSRS, M3FD, and RoadScene datasets. Bold and underlined values indicate the best and second-best results respectively. Tra.P. = Trainable Parameters, Fix.P. = Fixed Parameters.
@@
IV-C Comparison on Medical Datasets
To demonstrate the generalization capability of our method, we evaluate its performance on medical image fusion tasks. As shown in Table II, both our teacher and student networks achieve superior performance across different modality pairs. The improvement is particularly significant in structural preservation metrics (SSIM and $Q^{AB/F}$), which is crucial for medical applications.
Figure 4: Comprehensive visual comparison of different image fusion methods on Harvard Medical Image Fusion Datasets (PET-MRI, CT-MRI, SPECT-MRI). For each set of results, from left to right: Modality 1 Input, Modality 2 Input, Text-IF Output, Ours-Teacher Output, Ours-Distilled Output.
For SPECT-MRI fusion, our method shows notable advantages in preserving functional information while maintaining anatomical details. In CT-MRI fusion, where structural alignment is critical, our approach achieves the highest SSIM scores, indicating better structural preservation. The PET-MRI results further confirm our method’s effectiveness in handling multi-modal medical images with different characteristics. Fig. 4 presents visual examples from different medical modalities. Our fusion results exhibit better detail preservation and contrast enhancement, which is essential for clinical applications. The distilled student network maintains these advantages while significantly reducing computational requirements, making it more practical for clinical deployment.
TABLE II: Quantitative comparison with SOTA methods on medical image fusion tasks. Bold and underlined values indicate the best and second-best results respectively. The additional parameters (in parentheses) are non-trainable but used during inference.
@@ IV-C Comparison on Medical Datasets
IV-D Ablation Study
Figure 5: Overview of our ablation study.
IV-D1 Analysis of Teacher Network
Effect of Text Guidance.
To validate the effectiveness of text guidance, we compare our full model with its variant without text information. Removing text guidance leads to a significant performance drop across all metrics (EN: -0.023, VIF: -0.083, $Q^{AB/F}$: -0.05). This demonstrates that text descriptions provide valuable semantic information that helps the model better understand and fuse image features. The text guidance acts as an advanced form of reference information, offering rich contextual information beyond pixel-level features.
Impact of Different Loss Components.
We evaluate the contribution of each loss component by systematically removing them from the full model. The results, as shown in Table III, indicate that all loss components positively impact the final performance, although their importance varies. The intensity loss ($\mathcal{L}_{int}$) and gradient loss ($\mathcal{L}_{grad}$) have the most significant influence, highlighting their essential role in preserving structural information. The SSIM loss ($\mathcal{L}_{ssim}$) aids in maintaining perceptual quality, while the color loss ($\mathcal{L}_{color}$) ensures the natural appearance of fused results.
TABLE III: Ablation study on loss components for the teacher network. Results are reported on the MSRS test set.
[…]
$\mathcal{L}_{int}$ 6.695 1.027 0.697
$\mathcal{L}_{int}+\mathcal{L}_{color}$ 6.696 1.025 0.698
$\mathcal{L}_{int}+\mathcal{L}_{grad}+\mathcal{L}_{ssim}$ 6.731 1.058 0.730
$\mathcal{L}_{int}+\mathcal{L}_{color}+\mathcal{L}_{grad}$ 6.733 1.059 0.733
Feature Fusion Methods.
We investigate different fusion embedding strategies to combine features from infrared and visible images. The comparison includes concatenation & 1$\times$1 conv, unlearnable weighted attention, dynamic weighted conv, multiscale conv, and our proposed adaptive fusion method. As shown in Table IV, our adaptive fusion approach achieves superior performance by dynamically adjusting the fusion weights based on spatial and channel features.
TABLE IV: Comparison of different fusion modules. Results on MSRS test set.
[…]
Cat & 1$\times$1 conv 6.738 1.059 0.728
Unlearn weight atten 6.667 0.971 0.655
Dynamic weight conv 6.737 1.058 0.729
Multiscale conv 6.743 1.053 0.720
Ours 6.749 1.060 0.732
IV-D2 Analysis of Student Network
Effect of Model Size.
A notable finding in our distillation experiments is that reducing the model size by up to 90% (from 40.3M to 4.1M parameters if not calculate CLIP) does not significantly impact performance. As shown in Table V, the compact student model maintains comparable or even slightly better results across different metrics. This suggests that the original model might be over-parameterized for the current fusion task, and the knowledge distillation process effectively transfers essential fusion capabilities to a much smaller architecture.
TABLE V: Impact of model size on distilled network. Results on MSRS test set.
[…]
48-dim 39.1 6.780 1.071 0.734
32-dim 15.8 6.768 1.071 0.734
16-dim 4.1 6.763 1.075 0.734
Impact of Distillation Losses.
We evaluate the contribution of different components in our distillation framework. The results in Table VI show that result-level supervision (matching the teacher’s output) provides strong guidance, significantly improving the student model’s performance. The feature distillation loss, on the other hand, plays a key role in preserving feature-level consistency, ensuring that the student model captures important intermediate representations from the teacher. By introducing feature distillation, the model learns high-frequency details and intermediate features, enriching the information. This improves metrics like $EN$ and $Q^{AB/F}$, though it may slightly reduce the $VIF$ score due to imperfect alignment with perceptual fidelity.
TABLE VI: Ablation study on distillation losses (MSRS dataset).
[…]
$\mathcal{L}_{base}$ 6.728 0.989 0.681
$\mathcal{L}_{base}+\mathcal{L}_{res}$ 6.761 1.082 0.732
$\mathcal{L}_{base}+\mathcal{L}_{feat}+\mathcal{L}_{res}$ 6.763 1.075 0.734
Text Prior Analysis.
To better understand how textual priors influence the fusion process and how they are distilled, we conducted a detailed analysis of their effects at different network levels. As shown in Fig. 6, text guidance significantly impacts the model’s attention to different image regions. For instance, when processing a low-contrast nighttime scene, the text prior helps the model better distinguish the vehicle’s contours from the background, evident in the feature maps showing more defined object boundaries after prompt guidance. Different text descriptions (e.g., “Low Contrast” vs. “Low Light”) lead to distinct feature emphasis patterns. The distilled results show that our student network successfully inherits these text-guided attention patterns without explicit textual input, validating the effectiveness of our prior distillation approach in preserving semantic understanding.
Figure 6: Effect of text guidance on level 1 and fused image. (a) Original visible and infrared input images. (b) Feature maps before text guidance. (c) Feature maps after text guidance. (d) Final fused images under different text descriptions (“Low Contrast” and “Low Light”). (e) Distilled student network’s output.
Knowledge Distillation Analysis.
An intriguing observation from our experiments is that the distilled student network occasionally outperforms its teacher model. This phenomenon, illustrated in Fig. 7, suggests that the distillation process can act as a form of regularization. When training the teacher network with an extremely high weight for $\mathcal{L}_{ssim}$ (10$\times$ the baseline), the teacher becomes highly specialized in edge detection, sacrificing color fidelity. Surprisingly, the student, when distilled from such a teacher, learns to balance these competing objectives, maintaining edge clarity while preserving natural color representation. This indicates that extreme specialization in the teacher can provide a clearer supervision signal for certain features, enabling the student to learn more comprehensive fusion strategies. In our final implementation, a moderately increased weight for $\mathcal{L}_{ssim}$ proves sufficient for the student to achieve superior performance across multiple metrics with stable training.
Figure 7: When teacher performs extremely, the student learns more information. This figure illustrates how a student network learns to balance structural and color information even when the teacher over-prioritizes structural details, leading to color breakdown in the teacher’s output.
Our ablation studies reveal several key insights:
IV-E Inference Time Analysis
To evaluate the computational efficiency, we analyze the inference times of different components across the previous state-of-the-art (SoTA) text-guided method Text-IF, our teacher network, and the distilled student network. For text generation, we use Qwen2-VL [21] as the large language model (LLM), which operates on two NVIDIA RTX 4090 GPUs with the fast attention library [36]. Other timing experiments were conducted on a separate single NVIDIA RTX 4090 GPU, running with a batch size of 1 and an image resolution of $128\times 128$ pixels. Since these results may vary across different machines or even different states of the same machine, this comparison should be considered non-rigorous.
As shown in Table VII, the LLM component dominates the computational cost, requiring over 2 seconds per image. CLIP encoding adds another 110-115ms overhead. While the teacher network’s fusion module takes 133.4ms, our distilled student network reduces this to just 46.1ms—a 65% reduction in fusion time. More importantly, by eliminating the need for LLM and CLIP during inference, our student network achieves a dramatic 98% reduction in total inference time (from 2556.91ms to 46.11ms).
TABLE VII: Detailed inference time comparison (in milliseconds) for different components. LLM refers to the Qwen2-VL model used for text generation.
Method Data Load LLM CLIP Fusion Total
Text-IF 0.01 2261.2 115.7 152.1 2529.01
Teacher 0.01 2311.1 112.4 133.4 2556.91
Student 0.01 - - 46.1 46.11
Figure 8: Comprehensive visual comparison with different methods on various IVF datasets (MSRS, M3FD, RoadScene). For each set of results, from left to right: Visible Input, Infrared Input, Text-IF Output, Ours-Teacher Output, Ours-Distilled Output.
IV-F Limitations
TABLE VIII: Training time comparison for different epochs. All experiments were conducted on a single NVIDIA 4090 GPU.
Method 100 epochs 500 epochs
Text-IF 1.9h 11.1h
Teacher 2.0h 12.1h
Student (distill) 2.1h 13.5h
As shown in Table VIII, our distillation approach requires slightly longer training time compared to both the baseline Text-IF and the teacher model. For a 100-epoch training cycle, the student network’s distillation process takes 2.1 hours, approximately 10.5% longer than Text-IF (1.9h) and 5% longer than the teacher model (2.0h). This pattern persists for longer training durations, with 500-epoch training requiring 13.5 hours for distillation compared to 11.1 hours for Text-IF and 12.1 hours for the teacher model.
V Conclusion
We address the weak semantic understanding of traditional image fusion methods and the high cost of text-guided approaches by distilling textual priors and eliminating the need for text guidance during inference. Our approach, which leverages a teacher-student architecture and tailored prior distillation, significantly reduces the model size while retaining high performance. Extensive experiments and ablation studies validate the effectiveness of our method, highlighting its ability to achieve a strong trade-off between computational efficiency and fusion quality.